# CoSpaDi: Compressing LLMS via Calibration-Guided Sparse Dictionary Learning

### v2

@@ 3.2 Sparse Dictionary Learning
where $\mathbf{s}_{j}$ denotes the $j$-th column of the sparse coefficient matrix $\mathbf{S}$, and $\left\|\cdot\right\|_{0}$ denotes the $\ell_{0}$ pseudo-norm, which counts the number of nonzero entries in a vector.
The optimization problem in Eq. (4) is NP-hard in its original form and admits no closed-form solution. Nevertheless, it has been extensively studied, and several efficient algorithms have been proposed to obtain high-quality approximate solutions. Among these, K-SVD Aharon et al. (2006) is a well-established algorithm widely adopted across diverse applications.
K-SVD alternates between two main steps: (a) Sparse coding: Each signal $\mathbf{w}_{j}$ is approximated as a sparse linear combination of dictionary atoms (typically via orthogonal matching pursuit or LASSO). (b) Dictionary update: Each atom $\mathbf{d}_{i}$ is refined to better fit the data while preserving the sparsity pattern in $\mathbf{S}$. This is performed sequentially per atom: the algorithm identifies the subset of signals that activate $\mathbf{d}_{i}$, computes the residual error over those signals, and applies a rank-1 SVD to jointly update the atom and its associated coefficients.
3.3 CoSpadi: Activation-Aware SDL
In several recent works, including Chen et al. (2021), it has been observed that LLM weight matrices generally do not exhibit intrinsic low-rank structure. Consequently, direct low-rank approximation of the weights often leads to significant performance degradation. However, under the hypothesis that the input latent features to a given layer reside in a low-dimensional subspace, it remains feasible and often effective to approximate the corresponding layer weights as low-rank. This approach is well-motivated: the goal is not to preserve the weights in isolation, but rather to maintain the fidelity of the activations they produce when applied to low-dimensional inputs.
Within this framework, an effective approximation of $\mathbf{W}$ can be obtained by minimizing the output error:
$\tilde{\mathbf{W}}^{*}=\operatorname*{arg\,min}_{\tilde{\mathbf{W}}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\tilde{\mathbf{W}}\right\|_{F}\quad\text{s.t.}\quad\operatorname*{rank}(\mathbf{X}\tilde{\mathbf{W}})=r,$ (5)
where $\mathbf{X}\in\mathbb{R}^{N\times d_{1}}$ is a matrix of $N$ calibration input vectors $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d_{1}}\right\}_{i=1}^{N}$. This minimization admits the analytical solution: $\tilde{\mathbf{W}}=\mathbf{L}^{-1}\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\mathsf{T}}=\mathbf{B}\mathbf{C}$, where $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ is a non-singular transformation matrix derived from the calibration data (see Appendix A.2 for the derivation).
Motivated by this data-aware perspective, we propose adapting our sparse dictionary learning framework accordingly. Rather than directly approximating $\mathbf{W}$ as $\tilde{\mathbf{W}}=\mathbf{D}\mathbf{S}$, we instead minimize the reconstruction error of the output activations $\mathbf{Z}=\mathbf{X}\mathbf{W}$:
$\mathbf{D}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D},\mathbf{S}}\left\|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{D}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2}.$ (6)
To simplify this optimization, let $\mathbf{Y}=\mathbf{X}\mathbf{L}^{-1}\in\mathbb{R}^{N\times d_{1}}$ be a column-orthogonal matrix, i.e., $\mathbf{Y}^{\mathsf{T}}\mathbf{Y}=\mathbf{I}_{d_{1}}$, obtained by applying a linear transformation $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ to $\mathbf{X}$. We assume $N\geq d_{1}$ and that $\mathbf{X}$ has full column rank — conditions typically satisfied with sufficient calibration data. The matrix $\mathbf{L}$ can be computed via QR decomposition, SVD of $\mathbf{X}$, or Cholesky/eigen-decomposition of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
Introducing the auxiliary variables $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}$ and $\mathbf{D}_{L}=\mathbf{L}\mathbf{D}$, and noting that the Frobenius norm is invariant under left-multiplication by a column-orthogonal matrix, we can reformulate Eq. (6) as:
$\mathbf{D}_{L}^{\star},\mathbf{S}^{\star}=\operatorname*{arg\,min}_{\mathbf{D}_{L},\mathbf{S}}\left\|\mathbf{W}_{L}-\mathbf{D}_{L}\mathbf{S}\right\|_{F}^{2}\quad\text{s.t.}\quad\left\|\mathbf{s}_{j}\right\|_{0}\leq s\quad\forall\,j=1,\ldots,d_{2}.$ (7)
Since $\mathbf{L}$ is invertible, the final structured approximation of the original weights is given by:
$\tilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S},\quad\text{where}\quad\mathbf{D}_{a}=\mathbf{L}^{-1}\mathbf{D}_{L}$ (8)
is the activation-aware learned dictionary. Pseudocode for the full procedure is provided in Appendix A.3.
Cross-Layer SDL
Although most compression techniques focus on intra-layer optimizations, the repetitive layered structure of LLMs suggests the presence of significant inter-layer redundancy. To exploit this, we propose sharing a single dictionary across weight matrices of the same type from multiple layers. Specifically, let $\mathcal{G}$ denote a group of layer indices (e.g., all attention projection layers or all FFN layers). We form a grouped weight matrix by horizontally concatenating the corresponding layer weights: $\mathbf{W}_{G}=\begin{bmatrix}\mathbf{W}_{\ell_{1}},&\mathbf{W}_{\ell_{2}},&\ldots,&\mathbf{W}_{\ell_{L}}\end{bmatrix}\in\mathbb{R}^{d_{1}\times(d_{2}\cdot L)},$ where each $\mathbf{W}_{\ell}\in\mathbb{R}^{d_{1}\times d_{2}}$, and $L=|\mathcal{G}|$ is the number of layers in the group. We then apply the same activation-aware dictionary learning procedure as before, but now to $\mathbf{W}_{G}$. To ensure the approximation remains data-aware across all layers in the group, we construct a grouped calibration matrix by vertically stacking the corresponding input batches: $\mathbf{X}_{G}=\begin{bmatrix}\mathbf{X}_{\ell_{1}};&\mathbf{X}_{\ell_{2}};&\ldots;&\mathbf{X}_{\ell_{L}}\end{bmatrix}\in\mathbb{R}^{(N\cdot L)\times d_{1}}$ where each $\mathbf{X}_{\ell}\in\mathbb{R}^{N\times d_{1}}$ is the calibration input for layer $\ell$. The transformation matrix $\mathbf{L}\in\mathbb{R}^{d_{1}\times d_{1}}$ is then computed from $\mathbf{X}_{G}$ (e.g., via QR or Cholesky decomposition), ensuring that $\mathbf{Y}_{G}=\mathbf{X}_{G}\mathbf{L}^{-1}$ is column-orthogonal. The optimization proceeds by solving Eq. (7) with $\mathbf{W}_{L}=\mathbf{L}\mathbf{W}_{G}$, yielding a shared dictionary $\mathbf{D}_{a}$ and a sparse coefficient matrix $\mathbf{S}_{G}\in\mathbb{R}^{k\times(d_{2}\cdot L)}$.
@@ 3.3 CoSpadi: Activation-Aware SDL
This approach, inspired by Wang et al. (2024), reduces memory overhead by amortizing the dictionary cost across multiple layers, while preserving activation fidelity through data-aware calibration. Each layer’s compressed weights are then recovered by slicing the corresponding block from $\tilde{\mathbf{W}}_{G}=\mathbf{D}_{a}\mathbf{S}_{G}$.
Compression ratio.
For the low-rank setting, the compression ratio is $\gamma^{\mathrm{LR}}\coloneqq 1-\tfrac{r(d_{1}+d_{2})}{d_{1}d_{2}}$, where $r$ is the retained rank. For CoSpaDi, we store a dictionary and sparse codes plus a binary mask (details are provided in section A.4), giving
@@
$\gamma^{\mathrm{SD}}\coloneqq 1-\tfrac{\;\overbrace{d_{1}k}^{\text{dict. (bf16)}}+\overbrace{sd_{2}}^{\text{codes (bf16)}}+\overbrace{(kd_{2})/16}^{\text{mask (1 bit/entry)}}\;}{d_{1}d_{2}}.$ (9)
@@
Unlike the low-rank case, $\gamma^{\mathrm{SD}}$ depends on two knobs — the number of atoms $k$ and sparsity $s$ — allowing us to trade off model capacity and storage at fixed budget. We parameterize this trade-off by the ratio $\rho\coloneqq k/s$, which uniquely determines $(k,s)$ at a target $\gamma^{\mathrm{SD}}$:
@@
Inference Complexity
In terms of computational efficiency, low-rank and dictionary learning exhibit distinct inference-time complexity profiles: the former has a cost of $Tr(d_{1}+d_{2})$, whereas the latter – when exploiting sparsity and reusing inner products – achieves $Td_{1}K_{\text{active}}+Tsd_{2}$ (see section A.5), potentially yielding superior efficiency under favorable sparsity patterns. Although both methods share identical theoretical complexity under matched compression ratios, practical inference latency varies significantly due to factors such as the number of active atoms, indexing overhead, and hardware-specific kernel efficiency. Further details regarding complexity derivations are provided in Appendix A.5.
@@ 4 Experiments
In this section, we present our experimental setup. We first compare CoSpaDi with low-rank baselines in the per-layer setting, where each linear weight matrix is compressed independently. We then investigate cross-layer sharing, analyze the impact of coefficient matrix quantization, and conduct ablation studies on the $ks$-ratio.
@@ 4.1 Experimental Setup
We evaluate our method in both per-layer and cross-layer settings. For per-layer evaluations, we consider LLaMA-3.2 1B, Qwen-3 0.6B, LLaMA-3 8B, and Qwen-3 8B. For the cross-layer study, we follow the Basis Sharing Wang et al. (2024) setup and conduct experiments on LLaMA-2 7B. All models are evaluated in a zero-shot setting on PIQA Bisk et al. (2019), HellaSwag Zellers et al. (2019), OpenAI LAMBADA Paperno et al. (2016), ARC-easy and ARC-challenge Clark et al. (2018), SciQ Welbl et al. (2017), Race Lai et al. (2017), and MMLU Hendrycks et al. (2021), while perplexity is additionally reported on WikiText Merity et al. (2016) and LAMBADA-OpenAI. Compression is applied with target ratios of 0.2-0.5 with 0.1 step. For methods requiring calibration data, we randomly sample 256 sequences from the RefinedWeb dataset Penedo et al. (2023), each consisting of 1024 tokens. Unless specified otherwise, we compress all dense linear projections in self-attention (Q/K/V/O) and feed-forward network (gate/up/down). Embeddings and lm_head are left intact.
For low-rank methods the rank is uniquely defined with $\gamma^{LR}$ and we floored it to the nearest integer. For CoSpaDi we fixed $\rho=2$ ($k/s$ ratio) for all experiments, so $k$ and $s$ were obtained according to Eq. (10) and then were floored to the nearest integers.
In CoSpaDi we employ K-SVD using power iterations instead of full-svd for the dictionary updates. Specifically, we use 60 K-SVD iterations and 8 power iterations to ensure stable convergence. Further implementation details and convergence analysis are provided in Appendix A.8.
4.2 Ablation Study
Influence of $k/s$-ratio.
In the proposed CoSpaDi we can redistribute capacity across both $k$ and $s$ while preserving predefined compression ratio by varying the $\rho$ value ($k/s$). We verified its influence on the Llama 3.2 1B and report the metrics in Fig. 2.

### v3

@@ 3.3 Sparse dictionary learning: union of subspaces
where $\mathbf{s}_{j}$ is the $j$-th column of $\mathbf{S}$, $\left\|\cdot\right\|_{0}$ denotes the $\ell_{0}$ pseudo-norm, and $s$ indicates the sparsity level.
Compared to low-rank factorization, Eq. (5) differs in two crucial ways: (i) coefficients are sparse, so each column activates only $s$ atoms, and (ii) no orthogonality constraint is imposed on $\mathbf{D}$, allowing more flexible representations. As a result, columns can live in different subspaces spanned by different subsets of atoms: $\mathbf{W}$ is represented using a union of subspaces model rather than a single shared subspace.
3.4 CoSpaDi: activation-aware sparse dictionary learning for LLM projections
Motivated by the observation that activations, not projection weights, exhibit low-rank structure, we apply dictionary learning to LLM projection matrices under the activation-space objective (2). Concretely, we choose the constraint set $\mathcal{C}=\{\mathbf{D}\mathbf{S}:\mathbf{D}\in\mathbb{R}^{d_{1}\times k},\ \mathbf{S}\in\mathbb{R}^{k\times d_{2}},\ \|\mathbf{s}_{j}\|_{0}\leq s\}$ and solve
$\min_{\mathbf{D},\mathbf{S}}\ \|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{D}\mathbf{S}\|_{F}^{2}\quad\text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s,\ \forall j.$ (6)
Gram orthonormalization transform.
To simplify Eq. (6), consider the Gram matrix $\mathbf{G}\coloneqq\mathbf{X}^{\top}\mathbf{X}\in\mathbb{R}^{d_{1}\times d_{1}}$. We compute a Gram orthonormalization matrix $\mathbf{L}$ from the (uncentered) second moment $\mathbf{G}$ such that
$\mathbf{Y}\coloneqq\mathbf{X}\mathbf{L}^{-1}\quad\text{satisfies}\quad\mathbf{Y}^{\top}\mathbf{Y}=\mathbf{I},$ (7)
i.e., $\mathbf{Y}$ has orthogonal columns. In practice, $\mathbf{L}$ can be obtained via QR factorization of $\mathbf{X}$, or from a factorization of $\mathbf{G}$ (e.g., Cholesky when $\mathbf{G}$ is symmetric positive definite, and an SVD/eigendecomposition-based otherwise). Next, we define the transformed variables $\mathbf{W}_{L}\coloneqq\mathbf{L}\mathbf{W}$ and $\mathbf{D}_{L}\coloneqq\mathbf{L}\mathbf{D}$. Using the invariance of the Frobenius norm under left multiplication by a column-orthogonal matrix, Eq. (6) reduces to
$\min_{\mathbf{D}_{L},\mathbf{S}}\ \|\mathbf{W}_{L}-\mathbf{D}_{L}\mathbf{S}\|_{F}^{2}\quad\text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s,\ \forall j.$ (8)
After solving (8), we map back to the original parameterization:
$\tilde{\mathbf{W}}=\mathbf{D}_{a}\mathbf{S},\qquad\mathbf{D}_{a}=\mathbf{L}^{-1}\mathbf{D}_{L}.$ (9)
Thus, calibration data enters only through the “whitening” transform $\mathbf{L}$, while the core optimization reduces to standard sparse dictionary learning on $\mathbf{W}_{L}$.
Optimization via alternating minimization.
We approximately solve (8) by alternating between sparse coding and dictionary update:
  • Sparse coding. For each column $\mathbf{w}_{L,j}$ of $\mathbf{W}_{L}$, we compute a sparse code $\mathbf{s}_{j}$ with at most $s$ nonzeros by approximately solving $\min_{\mathbf{s}_{j}}\|\mathbf{w}_{L,j}-\mathbf{D}_{L}\mathbf{s}_{j}\|_{2}^{2}\ \text{s.t.}\ \|\mathbf{s}_{j}\|_{0}\leq s$, using orthogonal matching pursuit (OMP) (Tropp & Gilbert, 2007).
  • Dictionary update. With sparse supports fixed, we update the dictionary. Two common options are: (i) MOD (method of optimal directions), which updates all atoms jointly via a global least-squares step (Engan et al., 1999), and (ii) K-SVD, which updates atoms sequentially by solving a rank-1 approximation on a restricted residual (Aharon et al., 2006). Concretely, for atom $i$ let $\Omega_{i}=\{j:\ (\mathbf{s}_{j})_{i}\neq 0\}$ be the set of columns using atom $i$ and form the restricted residual $\mathbf{E}_{i}=\big(\mathbf{W}_{L}-\sum_{p\neq i}\mathbf{d}_{p}\mathbf{s}_{p,:}\big)_{[:,\Omega_{i}]}$. K-SVD updates $\mathbf{d}_{i}$ and $\mathbf{s}_{i,\Omega_{i}}$ via the top rank-1 component of $\mathbf{E}_{i}$, which can be computed exactly by SVD or approximated with truncated/randomized low-rank routines (Halko et al., 2011) or power iterations (Golub & Van Loan, 2013). In practice, both steps admit efficient implementations, e.g., Batch-OMP for sparse coding and approximate rank-1 updates for K-SVD, substantially reducing runtime and memory overhead (Rubinstein et al., 2008).
Across our ablations, K-SVD with power-iteration rank-1 updates as well as Batch-OMP provides the strongest performance under a reasonable wall-clock budget. The full procedure is summarized in Appendix 1.
3.5 Grouped / shared-dictionary extension
To exploit inter-layer redundancy, we optionally share a dictionary across a group of projections of the same type. Let $\mathcal{G}=\{\ell_{1},\ldots,\ell_{L}\}$ be a set of layers with weights $\mathbf{W}_{\ell}\in\mathbb{R}^{d_{1}\times d_{2}}$. We form a grouped matrix by concatenation
$\mathbf{W}_{\mathcal{G}}=\big[\mathbf{W}_{\ell_{1}},\ldots,\mathbf{W}_{\ell_{L}}\big]\in\mathbb{R}^{d_{1}\times(Ld_{2})}.$ (10)
To preserve functional fidelity across the group, we stack calibration inputs $\mathbf{X}_{\mathcal{G}}=\big[\mathbf{X}_{\ell_{1}}^{\top},\ldots,\mathbf{X}_{\ell_{L}}^{\top}\big]^{\top}\in\mathbb{R}^{(LN)\times d_{1}}$, compute a Gram orthonormalization matrix $\mathbf{L}$ from $\mathbf{X}_{\mathcal{G}}$, and solve (8) on $\mathbf{W}_{\mathcal{G},L}=\mathbf{L}\mathbf{W}_{\mathcal{G}}$ to obtain a shared dictionary and block-structured sparse codes. Each layer is then recovered by slicing the corresponding coefficient block.
@@ 3.5 Grouped / shared-dictionary extension
This approach, inspired by Basis Sharing (Wang et al., 2025a), reduces memory overhead by amortizing the dictionary cost across multiple layers, while preserving activation fidelity through data-aware calibration. Each layer’s compressed weights are then recovered by slicing the corresponding coefficient block. Specifically, if $\mathbf{S}_{\mathcal{G}}\in\mathbb{R}^{k\times(Ld_{2})}$ are the learned codes, then the codes for layer $\ell_{t}$ are $\mathbf{S}_{\ell_{t}}=\mathbf{S}_{\mathcal{G}}[:,(t-1)d_{2}:td_{2}]$.
3.6 Compression ratio
For low-rank factorization with rank $r$, storing two dense factors costs $r(d_{1}+d_{2})$ parameters, leading to a compression ratio:
$\gamma^{\mathrm{LR}}\coloneqq 1-\frac{r(d_{1}+d_{2})}{d_{1}d_{2}}.$ (11)
For CoSpaDi, we store a dense dictionary $\mathbf{D}\in\mathbb{R}^{d_{1}\times k}$ and sparse coefficients $\mathbf{S}\in\mathbb{R}^{k\times d_{2}}$. With $s$ nonzeros per column, the number of stored coefficient values is $sd_{2}$ (plus indices/mask; see Appendix A.4 for exact accounting). Thus, we can write
@@ 3.6 Compression ratio
$\gamma^{\mathrm{SD}}\coloneqq 1-\frac{\,\overbrace{d_{1}k}^{\text{dict.}}+\overbrace{sd_{2}}^{\text{values}}+\overbrace{(kd_{2})/16}^{\text{mask}}\,}{d_{1}d_{2}}.$ (12)
@@
Unlike low-rank compression, $\gamma^{\mathrm{SD}}$ depends on two knobs $(k,s)$, enabling additional flexibility at a fixed storage budget. We parameterize the trade-off via $\rho\coloneqq k/s$, which uniquely determines $(k,s)$ at a target $\gamma^{\mathrm{SD}}$:
@@
3.7 Inference complexity
In terms of computational efficiency, low-rank and dictionary learning exhibit distinct inference-time complexity profiles: the former has a cost of $Nr(d_{1}+d_{2})$, whereas the latter – when exploiting sparsity and reusing inner products – achieves $Nd_{1}K_{\text{active}}+Nsd_{2}$, potentially yielding superior efficiency under favorable sparsity patterns. Although both methods share identical theoretical complexity under matched compression ratios (Appendix A.5), practical inference latency varies significantly due to factors such as the number of active atoms, indexing overhead, and hardware-specific kernel efficiency. Further details regarding complexity derivations are provided in Appendix A.5.
@@ 4 Experiments
We evaluate CoSpaDi in two regimes: per-layer compression, where each projection matrix is compressed independently, and grouped (cross-layer) compression, where a dictionary is shared across a set of layers of the same type. We first present ablations that isolate key design choices (dictionary capacity allocation, data-awareness, packing/quantization, and solver variants), and then report main results for both regimes.
@@ 4.1 Experimental Setup
Models and layers.
We evaluate per-layer compression on LLaMA-3.2-1B, Qwen-3-0.6B, LLaMA-3-8B, Qwen-3-8B and Qwen-3-14B. For SVD-LLM we used the original code-base [^1] https://github.com/AIoT-MLSys-Lab/SVD-LLM with only first step on compression without extra finetuning. For grouped compression, we follow the Basis Sharing [^2] https://github.com/TUDa-HWAI/Basis_Sharing protocol and report results on LLaMA-2-7B. Unless stated otherwise, we compress all dense linear projections in self-attention (Q/K/V/O) and gated MLP (up/down/gate), while leaving embeddings and the LM head intact.
Calibration data and metrics.
For calibration we randomly sample 256 sequences of length 1024 from RefinedWeb (Penedo et al., 2023). We report standard zero-shot accuracy (normalized when available) on a suite of benchmarks (PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), OpenAI LAMBADA (Paperno et al., 2016), ARC-easy/ARC-challenge (Clark et al., 2018), SciQ (Welbl et al., 2017), Race (Lai et al., 2017), MMLU (Hendrycks et al., 2021)) and perplexity on WikiText (Merity et al., 2017) and OpenAI LAMBADA. We used lm-evaluation-harness 0.4.8 (Gao et al., 2024) to ensure reproducibility.
Compression parameterization.
For low-rank baselines, the target compression ratio (CR) uniquely determines the rank $r$. For CoSpaDi, CR is achieved by choosing the dictionary size $k$ and sparsity $s$ (nonzeros per column), with $\rho\coloneqq k/s$ controlling how capacity is split between breadth (more atoms) and depth (more active atoms per column).
Baselines.
In the per-layer setting we compare primarily to SVD-LLM (data-aware) (Wang et al., 2025c) and, where relevant, to plain truncated SVD (data-free). We also include comparisons to training-free structured pruning baselines.
We intentionally omit SVD-LLM v2 (Wang et al., 2025b) and Dobi-SVD (Qinsi et al., 2025) from the main baseline set for scope and fairness. Both methods emphasize dynamic (layer-wise) compression allocation, which is largely orthogonal to our contribution. Such allocation strategies are not specific to SVD and could be transferred to CoSpaDi by choosing $(k_{\ell},s_{\ell})$ per layer under a global budget. In addition, SVD-LLM v2 does not provide an official implementation, making it difficult to reproduce calibration and allocation details reliably and to apply them consistently in a dictionary-learning setting. Dobi-SVD further departs from our training-free protocol by requiring backpropagation to optimize the allocation. Its reported gains are also coupled to a quantization-specific “remapping” representation that changes the effective storage accounting. In particular, at moderate compression levels (e.g., CR <0.5<0.5), packing two 16-bit values into a single unit can lead to an effective overparameterization relative to standard bf16 storage. For these reasons, we focus on reproducible training-free baselines with standard storage conventions, and view dynamic allocation and quantization-specific representations as complementary extensions.
4.2 Ablation Studies
4.2.1 Capacity Allocation: the $k/s$ Ratio
At a fixed compression ratio (CR), CoSpaDi admits a family of factorizations parameterized by $\rho\coloneqq k/s$. Smaller $\rho$ increases per-column expressiveness (larger $s$) but reduces the dictionary size $k$; larger $\rho$ increases $k$ but restricts each column to fewer active atoms. We perform a coarse sweep over $\rho\in[1;5]$. Figure 2 (LLaMA-3.2-1B) shows that $\rho=2$ consistently provides the best trade-off across all considered CRs in terms of both average accuracy and perplexity. We therefore fix $\rho=2$ in all subsequent experiments for simplicity and comparability across models.
Notably, $\rho=1$ corresponds to the degenerate regime $k=s$, where each column may activate all atoms. In this case, the sparsity constraint becomes inactive and CoSpaDi reduces to a dense two-factor approximation, closely mirroring the “basis times coefficients” view underlying truncated SVD/PCA. This emphasizes that CoSpaDi can be seen as a strict generalization: choosing $\rho>1$ enforces sparsity and yields a union-of-subspaces model.